OpenStudy (anonymous):

if 40=44, 55=61. what does 47=?

OpenStudy (kinggeorge):

42. It's always 42.

OpenStudy (anonymous):

why?

OpenStudy (kinggeorge):

42 is the answer to life, the universe, and everything.

OpenStudy (anonymous):

and the oiptions are 49,50,51,52

OpenStudy (kinggeorge):

:(

OpenStudy (thomas5267):

@KingGeorge Good Job...

OpenStudy (anonymous):

:D

OpenStudy (thomas5267):

I have given the medal to you. The answer is soooo EPIC! @KingGeorge

OpenStudy (anonymous):

@thomas5267: Please don't take the following in the wrong way. There is nothing epochal about it. It's a cliched pun used extensively in M.SE and here.

OpenStudy (kinggeorge):

Definitely a cliched pun, but it's still fun to use every now and then. As for the actual solution, I might need a few minutes to see the pattern.

OpenStudy (anonymous):

52?

OpenStudy (anonymous):

yes,it is the answer .. but what is the explanation ?

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

??? :)

OpenStudy (anonymous):

Well, I won't spoil the fun :)

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/42_(number)

OpenStudy (lgbasallote):

ahh @FoolForMath he always gets the answers...as long as they make sense...he cant solve "special arithmetic" though hahahah =))))

OpenStudy (anonymous):

40+4=44 55+6=61 47+5=52